OpenStudy (anonymous):

convert 2x-3y=6 into slope intercept form

OpenStudy (amistre64):

/ it all by 3 and -the x part

OpenStudy (amistre64):

2/3x -y = 2 -y = -2/3 x +2 ; and then times it all by -1 to clean it up y = 2/3 x -2